Mark is 3 years older than Jane. In 4 years, Mark will be 1.5 times as old as Jane. How old is Mark now?
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

M=J+3
M+4=1.5(J+3)
J+3+4=1.5J+4.5
J+7=1.5J+4.5
J-1.5J=4.5-7
-.5J=-2.5
J=-2.5/-.5
J=5 YEARS OLD FOR jANE
M=5+3=8 YEARS OLD FOR MARK.
PROOF:
8+4=1.5(5+3)
12=1.5*8
12=12

Answer by Alan3354(69443)

Mark is 3 years older than Jane. In 4 years, Mark will be 1.5 times as old as Jane. How old is Mark now?
0 solutions
----------------
M = J + 3 (age difference now)
In 4 years,
M + 4 = 1.5*(J+4)
Substitute for M:
J+3+4 = 1.5*(J+4)
J+7 = 1.5J + 6
Multiply by 2
2J+14 = 3J + 12
J = 2
M = 5